OpenStudy (monkeyboom):

does anyones know how i can solve this 3x^2+10-8=0

OpenStudy (aaronandyson):

3x^2 + 10 - 8 = 0 3x^2 + 2 = 0 3x^2 = -2 Is there any other information given?

OpenStudy (monkeyboom):

no

OpenStudy (monkeyboom):

i think u have to factorise

OpenStudy (aaronandyson):

3x^2 + 10 - 8 = 0 3x^2 + 2 = 0 3x^2 = -2 \[x^2 = -\frac{ 2 }{ 3 }\]

OpenStudy (aaronandyson):

What are the option given to you?

OpenStudy (monkeyboom):

the answere in the book says x=-4 or x=2/3

OpenStudy (aaronandyson):

I think you've missed something while copying the question...please post a screenshot of the original question.

OpenStudy (aaronandyson):

Is this the original question: \[3x^2 + 10x - 8 = 0?\

OpenStudy (aaronandyson):

\[3x^2 + 10x - 8 = 0\]

OpenStudy (aaronandyson):

Factorise it.

OpenStudy (monkeyboom):

yes

OpenStudy (aaronandyson):

Use this method. 3x^2 + 10x - 8 = 0 Split the middle sum such that it's sum is equal to the product of the first term and the last term. i.e multiple 3x^2(the first digit ) with 8 (the second digit) So, 3x^2 multiplied by 8 = ?

OpenStudy (monkeyboom):

24

OpenStudy (aaronandyson):

Yes. 3 into 8 is 24. So 3x^2 into 8 = 24x^2,right? So now split the middle digit(that is 10) such that these two number when multiplies give 24x^2 as the product.

OpenStudy (monkeyboom):

5

OpenStudy (aaronandyson):

no no no.. 5 times 5 is 25. Those two numbers are 6x and 4x.

OpenStudy (monkeyboom):

oh ok

OpenStudy (aaronandyson):

3x^2 + 10x - 8 =0 Will give (x+4)(3x - 2) = 0

OpenStudy (aaronandyson):

x is -4 or x is 2/3

OpenStudy (monkeyboom):

oh omg thank u so much aaron!